# Asymmetric Perturbation in Solving Bilinear Saddle-Point Optimization

### v1

@@ 6 Related literature
As an alternative approach, payoff perturbation has gained renewed attention. In this approach, players’ payoff functions are regularized with strongly convex terms (Cen et al., 2021, 2023; Pattathil et al., 2023), which stabilizes the dynamics and leads to convergence. Some existing works have shown convergence to an approximate equilibrium under fixed perturbation (Sokota et al., 2023; Tuyls et al., 2006; Coucheney et al., 2015; Leslie and Collins, 2005; Abe et al., 2022; Hussain et al., 2023). To recover equilibria of the original game, later studies have employed a decreasing schedule or iterative regularization (Facchinei and Pang, 2003; Koshal et al., 2013; Yousefian et al., 2017; Bernasconi et al., 2022; Liu et al., 2023; Cai et al., 2023), or have updated the regularization center periodically (Perolat et al., 2021; Abe et al., 2023, 2024). In contrast to these approaches, our algorithms converge without decaying or modifying the perturbation.
Extensive-form games, which model sequential and imperfect-information interactions, have also been studied from both theoretical and empirical perspectives. Strategy representations can be broadly classified into sequence-form (von Stengel, 1996) and behavioral-form. Under the sequence-form, optimistic algorithms have been shown to enjoy last-iterate convergence guarantees (Lee et al., 2021). Under the behavioral-form, perturbation-based approaches have also been applied successfully (Perolat et al., 2021; Sokota et al., 2023; Liu et al., 2023). Our asymmetric perturbation is compatible with both representations. Notably, our theoretical results in Theorem 4.1 and Corollary 4.2 hold under the sequence-form, and we empirically demonstrate strong performance under the behavioral-form as well.
@@ 7 Conclusion
This paper introduces an asymmetric perturbation technique for solving saddle-point optimization problems, addressing key challenges in learning dynamics and equilibrium computation. Unlike symmetric perturbation methods that yield only approximate equilibria, our approach guarantees last-iterate convergence to the exact equilibrium without requiring parameter adjustment. This structural insight opens promising directions for future research, including the development of learning algorithms that further exploit asymmetric perturbation to enhance stability and convergence.
@@ A.2 Proximity to equilibrium under symmetric and asymmetric perturbations in Biased Matching Pennies
Figure 5 exhibits the proximity of $x^{\mu}$ to $x^{\ast}$ as $\mu$ varies. Notably, under the symmetric perturbation, $x^{\mu}$ coincides with $x^{\ast}$ when $\mu$ is set to $\mu=\frac{\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)-v^{\ast}}{\left\|x^{\ast}\right\|^{2}-\frac{1}{m}}=\frac{4}{3}$. This result underscores the statement in Theorem 3.1, that $x^{\mu}$ does not coincide with $x^{\ast}$ as long as $\mu\neq\frac{\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)-v^{\ast}}{\left\|x^{\ast}\right\|^{2}-\frac{1}{m}}$.
@@
Figure 5: The proximity of $x^{\mu}$ to $x^{\ast}$ under the symmetric perturbation and the asymmetric perturbation with varying $\mu$. The game matrix $A$ is given by $[[\frac{1}{3},-\frac{2}{3}],[-\frac{2}{3},1]]$.
@@ A.3 Additional experiments in normal-form games
Figure 7: Trajectories for SymP-GDA (top row) and AsymP-GDA (bottom row) under different perturbation strengths $\mu\in\{0.5,1.0,4.0\}$ in BRPS.
@@
Figure 7 illustrates the trajectories of SymP-GDA (top row) and AsymP-GDA (bottom row) under varying perturbation strengths $\mu\in\{0.5,1.0,2.0,4.0\}$ in BRPS. For SymP-GDA, the trajectories do not converge directly to the equilibrium even for small values of $\mu=0.5,1.0$. Instead, they follow circuitous and elongated paths, resulting in slower convergence. Conversely, as $\mu$ increases ($\mu=2.0,4.0$), the trajectories become more direct, leading to faster convergence, but it remains farther from the equilibrium. In contrast, AsymP-GDA leads to direct convergence to the equilibrium with small perturbation strengths. For $\mu$ values up to $2.0$ the trajectories converge directly to the equilibrium. However, as $\mu$ increases beyond a threshold ($\mu=4.0$), the trajectory deviates from the equilibrium. These results provide a more detailed understanding of the trends observed in Figures 1 and 1, further illustrating the differences in convergence dynamics between symmetric and asymmetric perturbations.
A.4 Pseudo Code for Experiments in Extensive-Form Games
In this section, we present the pseudocode for AsymP-CFR+ in Algorithm 1.
Algorithm 1 AsymP-CFR+
1:  $(\boldsymbol{R}^{x}_{t},\boldsymbol{R}^{y}_{t})=0$
2:  for $t=0$ to $T$ do
3:     $\boldsymbol{x}_{t}=\left[\boldsymbol{R}^{x}_{t}\right]^{+}\big{/}\|\left[\boldsymbol{R}^{x}_{t}\right]^{+}\|_{1}$
4:     $\boldsymbol{R}^{x}_{t+1}=\left[\boldsymbol{R}^{x}_{t}+\boldsymbol{r}({x}_{t})-\mu\boldsymbol{x}_{t}\right]^{+}$
5:     $\boldsymbol{y}_{t}=\left[\boldsymbol{R}^{y}_{t}\right]^{+}\big{/}\|\left[\boldsymbol{R}^{y}_{t}\right]^{+}\|_{1}$
6:     $\boldsymbol{R}^{y}_{t+1}=\left[\boldsymbol{R}^{y}_{t}+\boldsymbol{r}({y}_{t})\right]^{+}$
7:  end for
Appendix B Independently Perturbed Game
@@
Appendix C Adaptively Asymmetric Perturbation
As noted in Remark 4.3, AsymP-GDA may require a very small value of $\mu$ to ensure convergence to an equilibrium in the original game. To address this limitation, this section introduces an enhanced variant called Adaptively AsymP-GDA (Ada-AsymP-GDA), which incorporates an adaptive anchoring strategy $\sigma\in\mathcal{X}$ [Perolat et al., 2021, Abe et al., 2024].
In Ada-AsymP-GDA, instead of using the perturbation term $\mu x^{t}$ as in the original algorithm, we apply the gradient of the squared distance between $x^{t}$ and the anchor $\sigma$, namely $\mu(x^{t}-\sigma)$. The anchoring strategy $\sigma$ is updated periodically: it is reset to the current strategy $x^{t}$ every $T_{\sigma}$ iterations.
Let $k(t)$ denote the number of times $\sigma$ has been updated up to iteration $t$, and let $\sigma^{k(t)}$ denote the anchoring strategy after $k(t)$ updates. Since $\sigma$ is updated every $T_{\sigma}$ iterations, we have $k(t)=\lfloor(t-1)/T_{\sigma}\rfloor+1$ and $\sigma^{k(t)}=x^{T_{\sigma}(k(t)-1)+1}$. In summary, Ada-AsymP-GDA updates each player’s strategy at iteration $t\in[T]$ according to:
@@ Appendix C Adaptively Asymmetric Perturbation
C.1 Last-Iterate Convergence Rate
We now establish the last-iterate convergence rate of Ada-AsymP-GDA. In particular, we show that the final strategy $x^{T+1}$ converges to an equilibrium of the original game at a rate of $\mathcal{O}(1/T)$.
Theorem C.1.
Let $\mu>0$ be an arbitrary perturbation strength. If we set $\eta<\min\left(\frac{\mu}{2(\mu^{2}+\|A\|^{2})},\frac{8\left(\mu+\|A\|\right)}{D\min\left(\mu,\frac{\|A\|^{2}}{\mu}\right)}\right)$ and $T_{\sigma}=cT$ for some constant $c\leq\min\left(1,\frac{\alpha^{2}}{2\mu^{2}\left\|\prod_{\mathcal{X}^{\ast}}(x^{1})-x^{1}\right\|^{2}}\right)$, then the strategy $x^{T+1}$ satisfies:
$\left\|\prod_{\mathcal{X}^{\ast}}(x^{T+1})-x^{T+1}\right\|^{2}$ $\leq\mathcal{O}\left(\frac{1}{T}\right).$
Notably, Ada-AsymP-GDA achieves a last-iterate convergence rate of $\mathcal{O}(1/T)$ even when using a relatively large value of $\mu$.
C.2 Empirical performance in normal-form games
In this section, we empirically evaluate the performance of Ada-AsymP-GDA on the same normal-form games presented in Appendix A.3. We compare Ada-AsymP-GDA against AsymP-GDA, SymP-GDA, and Ada-SymP-GDA, where Ada-SymP-GDA applies perturbations $\mu(x^{t}-\sigma_{x}^{k(t)})$ and $\mu(y^{t}-\sigma_{y}^{k(t)})$ to both players’ gradients.
For all algorithms, we use a learning rate of $\eta=0.01$ and a perturbation strength of $\mu=5$. The update interval for the anchoring strategy is set to $T_{\sigma}=10,000$ for both Ada-AsymP-GDA and Ada-SymP-GDA.
Figure 8 illustrates the logarithm of NashConv averaged over $100$ random seeds. Ada-AsymP-GDA exhibits substantially faster convergence to an equilibrium compared to the other methods.
(a) BRPS
(b) M-Ne
Figure 8: Performance of Ada-AsymP-GDA in normal-form games. The shaded area represents the standard errors.
C.3 Empirical performance in extensive-form games
We also extend our method to extensive-form games through a variant called Adaptive AsymP-CFR+ (Ada-AsymP-CFR+), which integrates the adaptive anchoring strategy into AsymP-CFR+. Specifically, in Ada-AsymP-CFR+, the cumulative counterfactual regret $\boldsymbol{R}_{t}^{x}$ for player $x$ is updated as:
$\boldsymbol{R}^{x}_{t+1}=\left[\boldsymbol{R}^{x}_{t}+\boldsymbol{r}({x}_{t})-\mu(\boldsymbol{x}_{t}-\sigma^{k(t)})\right]^{+}.$
We compare Ada-AsymP-CFR+ with AsymP-CFR+, SymP-CFR+, and Ada-SymP-CFR+ in the same extensive-form games in Section 5. Note that Ada-SymP-CFR+ applies symmetric perturbations with the adaptive anchoring strategy. For all algorithms, we use a perturbation strength of $\mu=0.05$. The update interval for the anchoring strategy is set to $T_{\sigma}=2,500$ for both Ada-AsymP-CFR+ and Ada-SymP-CFR+.
Figure 9 illustrates the NashConv values for each game. The results indicate that Ada-AsymP-CFR+ consistently achieves lower NashConv values, demonstrating superior convergence performance.
Figure 9: Performance of Ada-AsymP-CFR+ in extensive-form games.
Appendix D Proof of Theorem 3.1

### v2

@@ 6 Related Literature
As an alternative approach, payoff perturbation has gained renewed attention. In this approach, players’ payoff functions are regularized with strongly convex terms (Cen et al., 2021, 2023; Pattathil et al., 2023), which stabilizes the dynamics and leads to convergence. Some existing works have shown convergence to an approximate equilibrium under fixed perturbation (Sokota et al., 2023; Tuyls et al., 2006; Coucheney et al., 2015; Leslie and Collins, 2005; Abe et al., 2022; Hussain et al., 2023). To recover equilibria of the original game, later studies have employed a decreasing schedule or iterative regularization (Facchinei and Pang, 2003; Koshal et al., 2013; Yousefian et al., 2017; Bernasconi et al., 2022; Liu et al., 2023; Cai et al., 2023), or have updated the regularization center periodically (Perolat et al., 2021; Abe et al., 2023, 2024). In contrast to these approaches, our algorithms avoid iteration-budget-dependent tuning of the perturbation strength by exploiting the equilibrium invariance property of our asymmetric perturbation.
Extensive-form games, which model sequential decision-making under imperfect information, have been studied extensively from both theoretical and empirical perspectives. Regarding last-iterate convergence, Lee et al. (2021) establish last-iterate convergence guarantees for optimistic algorithms in the sequence-form representation (von Stengel, 1996). More recently, Liu et al. (2023) show last-iterate convergence for symmetrically perturbed algorithms with a decreasing schedule on the perturbation strength. In contrast, our asymmetric perturbation avoids such delicate scheduling of the perturbation strength, and thus may be more amenable to large-scale extensive-form settings where optimistic algorithms can be difficult to deploy reliably.
@@ 7 Conclusion
This paper introduces an asymmetric perturbation technique for solving saddle-point optimization problems, addressing key challenges in learning dynamics and equilibrium computation. Unlike symmetric perturbation, which in general yields only approximate equilibria for any fixed perturbation strength, our approach admits an invariance regime: for sufficiently small $\mu$, the minimax strategy of the perturbed game coincides with that of the original game. Building on this structural property, we propose AsymP-GDA and establish a linear last-iterate convergence rate, improving over symmetric perturbation approaches in the same bilinear setting. Moreover, we show that the same invariance phenomenon can be leveraged to design a parameter-free procedure that retains a linear rate without requiring knowledge of game-dependent constants.
Our theoretical results target bilinear two-player zero-sum games. The key insight of equilibrium invariance in Theorem 3.1 relies on the near-linear growth of the objective function. We believe that an analogous formulation can be posed beyond bilinear games, including two-player zero-sum Markov games. Moreover, extending asymmetric perturbation to behavioral-form representation in extensive-form games and analysis of its performance are promising future directions.
@@ A.2 Proximity to Equilibrium under Symmetric and Asymmetric Perturbations in Biased Matching Pennies
Figure 5 exhibits the proximity of $x^{\mu}$ to $x^{\ast}$ as $\mu$ varies. Notably, under the symmetric perturbation, $x^{\mu}$ coincides with $x^{\ast}$ when $\mu$ is set to $\mu=\frac{\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)-v^{\ast}}{\left\|x^{\ast}\right\|^{2}-\frac{1}{m}}=\frac{4}{3}$. This result underscores the statement in Theorem B.1, that $x^{\mu}$ does not coincide with $x^{\ast}$ as long as $\mu\neq\frac{\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)-v^{\ast}}{\left\|x^{\ast}\right\|^{2}-\frac{1}{m}}$.
@@
Figure 5: The proximity of $x^{\mu}$ to $x^{\ast}$ under the symmetric perturbation and the asymmetric perturbation with varying $\mu$. The game matrix AA is given by $\frac{1}{3},-\frac{2}{3}],[-\frac{2}{3},1]]$.
@@ A.3 Additional Experiments in Normal-Form Games
Figure 7: Trajectories for SymP-GDA (top row) and AsymP-GDA (bottom row) under different perturbation strengths $\mu\in\{0.5,1.0,2.0,4.0\}$ in BRPS.
Figure 8: Sensitivity to the perturbation strength for AsymP-GDA with $\eta=0.01$ in BRPS and M-Ne.
@@
Figure 7 illustrates the trajectories of SymP-GDA (top row) and AsymP-GDA (bottom row) under varying perturbation strengths $\mu\in\{0.5,1.0,2.0,4.0\}$ in BRPS. For SymP-GDA, the trajectories do not converge directly to the equilibrium even for small values of $\mu=0.5,1.0$. Instead, they follow circuitous and elongated paths, resulting in slower convergence. Conversely, as $\mu$ increases ($\mu=2.0,4.0$), the trajectories become more direct, leading to faster convergence, but they remain farther from the equilibrium. In contrast, AsymP-GDA leads to direct convergence to the equilibrium with small perturbation strengths. For $\mu$ values up to $2.0$ the trajectories converge directly to the equilibrium. However, as $\mu$ increases beyond a threshold ($\mu=4.0$), the trajectory deviates from the equilibrium. These results provide a more detailed understanding of the trends observed in Figures 1 and 1, further illustrating the differences in convergence dynamics between symmetric and asymmetric perturbations.
Figure 8 illustrates the performance of AsymP-GDA on BRPS and M-Ne with $\mu\in\{0.01,0.1,1.0,2.0,4.0\}$ and $\eta=0.01$. For sufficiently small $\mu$, the limit point coincides with an equilibrium of the original game. However, decreasing $\mu$ also slows convergence. Overall, these results highlight a trade-off between accuracy and convergence speed.
A.4 Hyperparameter Settings for Extensive-Form Games
For the experiments in Section 5, we tuned the hyperparameters of each algorithm separately for each game to obtain the best performance. The resulting settings are summarized in Table 4. For the dilated regularizer, we use the unweighted version throughout, setting $\alpha_{i}=1$ for all $i$, as in Lee et al. [2021].
A.5 Comparison with CFR-Based Algorithms
This section compares the performance of AsymP-DGDA with several CFR-based algorithms, including CFR [Zinkevich et al., 2007], CFR+ [Tammelin, 2014], Discounted CFR (DCFR), and Linear CFR (LCFR) [Brown and Sandholm, 2019]. We use the same games as in Section 5 and keep the hyperparameter settings for AsymP-DGDA and SymP-DGDA unchanged. Figure 9 shows the logarithm of NashConv as a function of the number of strategy updates. For AsymP-DGDA and SymP-DGDA, we report NashConv for the last-iterate strategies. For the CFR-based algorithms, we report NashConv for the average-iterate strategies. Overall, AsymP-DGDA achieves lower NashConv than the CFR-based baselines on most games, with Leduc Poker being the exception.
Game Algorithm $\eta$ $\mu$
Kuhn Poker DMWU 0.01 -
DGDA 0.01 -
DOMWU 0.1 -
DOGDA 0.1 -
SymP-DGDA 0.1 0.001
AsymP-DGDA 0.1 0.01
Leduc Poker DMWU 0.05 -
DGDA 0.01 -
DOMWU 0.1 -
DOGDA 0.1 -
SymP-DGDA 0.05 0.0001
AsymP-DGDA 0.1 0.0001
Liars Dice (4 sides) DMWU 0.01 -
DGDA 0.01 -
DOMWU 0.1 -
DOGDA 0.1 -
SymP-DGDA 0.01 0.0001
AsymP-DGDA 0.1 0.001
Goofspiel (4 cards) DMWU 0.1 -
DGDA 0.1 -
DOMWU 0.01 -
DOGDA 0.01 -
SymP-DGDA 0.1 0.0001
AsymP-DGDA 0.1 0.05
Goofspiel (5 cards) DMWU 0.1 -
DGDA 0.1 -
DOMWU 0.01 -
DOGDA 0.01 -
SymP-DGDA 0.1 0.0001
AsymP-DGDA 0.1 0.001
Table 4: Hyperparameters
Figure 9: Performance in extensive-form games. AsymP-DGDA performs two strategy updates per iteration for each player. For a fair comparison across methods with different per-iteration computational costs, we report the total number of strategy updates on the x-axis rather than iterations.
Appendix B Impossibility Results for Symmetric Perturbation
As we stated in Section 2, existing works [Liu et al., 2023, Abe et al., 2024] have shown that the distance between the solution of the symmetrically perturbed game $(x^{\mu},y^{\mu})$ and the solution in the original game $(x^{\ast},y^{\ast})$ is upper bounded by $\mathcal{O}(\mu)$. However, they do not guarantee that the two solutions coincide, even for a small $\mu>0$. In contrast, the following theorem provides the first formal impossibility result, proving that $(x^{\mu},y^{\mu})$ almost never coincides with $(x^{\ast},y^{\ast})$.
Theorem B.1.
The proof is provided in Appendix H. Additionally, we extend our analysis to the case where both players have different perturbation strengths, i.e., $\mu_{x}>0$ and $\mu_{y}>0$, as shown in Appendix C.
Discussion on Theorem B.1.
When the game is symmetric.
Corollary B.2.
Assume that $A^{\top}=-A$. Under the same setup as Theorem B.1, the equilibrium $(x^{\mu},y^{\mu})$ in (2) always satisfies $x^{\mu}\neq x^{\ast}$ and $y^{\mu}\neq y^{\ast}$ for any $\mu>0$.
This is because it always holds that $v^{\ast}-\left(\frac{\mathbf{1}_{m}}{m}\right)^{\top}A\left(\frac{\mathbf{1}_{n}}{n}\right)=0$ when $A^{\top}=-A$. Figure 1 shows the proximity of $x^{\mu}$ to $x^{\ast}$ with varying perturbation strength $\mu$ in a simple biased Rock-Paper-Scissors game. We observe that as long as $\mu>0$, $x^{\mu}$ remains distant from $x^{\ast}$. This observation supports the theoretical results in Theorem B.1 and Corollary B.2.
Appendix C Independently Perturbed Game
@@
2 $(\hat{x}_{x}^{\mu_{0}},\hat{y}_{x}^{\mu_{0}})\leftarrow(x_{\mathrm{init}},y_{\mathrm{init}}),~(\hat{x}_{y}^{\mu_{0}},\hat{y}_{y}^{\mu_{0}})\leftarrow(x_{\mathrm{init}},y_{\mathrm{init}})$
3 for $k=1,2,\cdots$ do
4 $\mu_{k}\leftarrow\frac{\mu_{\mathrm{init}}}{2^{k-1}}$
5 $\eta_{k}\leftarrow\min\left(\eta_{k-1},\frac{\mu}{\mu^{2}+mn}\right)$
6 $\delta_{k}\leftarrow\frac{\mu_{k}}{2\|A\|^{2}\max_{z\in\mathcal{Z}}\|z\|^{2}}\varepsilon^{2}$
7 $(x_{x}^{1},y_{x}^{1})\leftarrow(\hat{x}_{x}^{\mu_{k-1}},\hat{y}_{x}^{\mu_{k-1}})$
8 $(x_{y}^{1},y_{y}^{1})\leftarrow(\hat{x}_{y}^{\mu_{k-1}},\hat{y}_{y}^{\mu_{k-1}})$
9 $(\hat{x}_{x}^{\mu_{k}},\hat{y}_{x}^{\mu_{k}})\leftarrow\textsc{AsymP-GDA}_{x}(x_{x}^{1},y_{x}^{1},\eta_{k},\mu_{k},\delta_{k})$
10 $(\hat{x}_{y}^{\mu_{k}},\hat{y}_{y}^{\mu_{k}})\leftarrow\textsc{AsymP-GDA}_{y}(x_{y}^{1},y_{y}^{1},\eta_{k},\mu_{k},\delta_{k})$
11 if $\max_{x\in\mathcal{X}}\langle A\hat{y}_{y}^{\mu_{k}},\hat{x}_{x}^{\mu_{k}}-x\rangle+\max_{y\in\mathcal{Y}}\langle-A^{\top}\hat{x}_{x}^{\mu_{k}},\hat{y}_{y}^{\mu_{k}}-y\rangle\leq\varepsilon$ then
12    return $\hat{x}_{x}^{\mu_{k}},\hat{y}_{y}^{\mu_{k}}$
13   end if
14
15 end for
-3mm
1mm
16 subroutine $\textsc{AsymP-GDA}_{x}(x^{1},y^{1},\eta,\mu,\delta)$
17 for $t=1,2,\cdots$ do
18    $x^{t+1}=\Pi_{\mathcal{X}}\left(x^{t}-\eta(Ay^{t}+\mu x^{t})\right)$
19    $y^{t+1}=\Pi_{\mathcal{Y}}\left(y^{t}+\eta A^{\top}x^{t+1}\right),$~
20    if $\max_{x\in\mathcal{X}}\langle Ay^{t+1}+\mu x^{t+1},x^{t+1}-x\rangle+\max_{y\in\mathcal{Y}}\langle-A^{\top}x^{t+1},y^{t+1}-y\rangle\leq\delta$ then
21       return $(x^{t+1},y^{t+1})$
22
23
24   end for
25
26
-3mm
1mm
27 subroutine $\textsc{AsymP-GDA}_{y}(x^{1},y^{1},\eta,\mu,\delta)$
28 for $t=1,2,\cdots$ do
29    $y^{t+1}=\Pi_{\mathcal{Y}}\left(y^{t}+\eta(A^{\top}x^{t}-\mu y^{t})\right),$~
30    $x^{t+1}=\Pi_{\mathcal{X}}\left(x^{t}-\eta Ay^{t+1}\right)$
31    if $\max_{x\in\mathcal{X}}\langle Ay^{t+1},x^{t+1}-x\rangle+\max_{y\in\mathcal{Y}}\langle-A^{\top}x^{t+1}+\mu y^{t+1},y^{t+1}-y\rangle\leq\delta$ then
32       return $(x^{t+1},y^{t+1})$
33
34
35   end for
36
37
Algorithm 1 Adaptive-$\mu$ AsymP-GDA
Appendix D Parameter-Free AsymP-GDA
As noted in Remark 3.2, AsymP-GDA may require a very small perturbation strength $\mu$ to guarantee convergence to an equilibrium of the original game. To address this limitation, this section introduces a parameter-free variant (Algorithm 1) that adaptively shrinks $\mu$.
Algorithm 1 takes a target NashConv value $\varepsilon$ as an input, and runs AsymP-GDA over multiple episodes. In each episode, the algorithm runs AsymP-GDA until the duality gap of the current strategies falls below $\Theta(\varepsilon^{2})$. It then checks whether the resulting strategy profile has a NashConv value at most $\varepsilon$. If the condition is satisfied, the algorithm terminates and outputs the strategies. Otherwise, it halves the perturbation strength $\mu$ and proceeds to the next episode.
D.1 Last-Iterate Convergence Rate
We establish the last-iterate convergence rate for Algorithm 1. Without loss of generality, we assume $\max_{x\in\mathcal{X}}\|A^{\top}x\|_{\infty}\leq 1$, $\max_{y\in\mathcal{Y}}\|Ay\|_{\infty}\leq 1$, and $\max_{(i,j)\in[m]\times[n]}|A_{ij}|\leq 1$ since these conditions can be satisfied by re-scaling $A$. Under this assumption, we show that the iteration complexity for the algorithm with an arbitrarily large initial perturbation strength is $\mathcal{O}(\ln(1/\varepsilon))$, in Theorem 4.3,
Theorem 4.3 implies that Algorithm 1 attains a linear last-iterate convergence rate, i.e., $\mathcal{O}(\exp(-t))$, even when the perturbation strength $\mu$ is initialized arbitrarily large. While Liu et al. [2023] have adopted a similar shrinking mechanism, their analysis yields a $\tilde{\mathcal{O}}(1/t)$ rate, which is slower than the linear rate established here.
Appendix E Proof of Theorem 3.1